Operation Summary for July 22, 2011 - July 25, 2011 |
The Integrated Luminosity for the period from 7/18/11 to 7/25/11 was 46.11 Pb-1 (inverse picobarns). NuMI remained off due to a magnet problem. Friday July 22 The day shift began with Operations monitoring the Tevatron (TEV) store 8912 (59.45E30) and the Antiproton Source (Pbar) stack (76.4E10), with no Recycler stash due to a Main Injector (MI) in access, with Linac, and Booster operating normally, with the MiniBooNE down due to the MI access, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, operators reported at 8:48 AM that a LCW leak was found on a quadrupole magnet (Q847) in MI-12B; Water Works personnel investigated. At 9:42 AM, the TEV store 8912 aborted due to a lightning strike; there was no quench. Operators prepared CDF for an access at 9:45 AM. Operators prepared the AP-3 transfer line for a reverse antiproton study at 10 AM. Operators prepared Booster for access at 10:16 AM. Operators put Pbar into standby mode and prepared the TEV for access. Operators prepared D0 for an access at 10:56 AM. Booster came out of access at 12:55 PM. D0 came out of access at 12:58 PM. EE Support personnel investigating the NuMI corrector magnet (V108) reported at 1:31 PM that six of the sixteen cables going to the magnet were shorted together. The work in MI-12B finished and operators began turning on MI and MiniBooNE at 3 PM. FESS personnel began cleaning the TEV sector A4 pond water strainer at 3:18 PM. Pbar began stacking at 5:48 PM. Operators reported that the following work was completed: TEV vacuum and Tevatron Electron Lens, MI V108 diagnosis and LCW leak repairs, MI-12 LCW leak repair, and a Booster RF power amplifier (BRF10) replacement. During the evening shift, Operations established beam to MiniBooNE at 4:43 PM. MTA began a one-hour study at 4:44 PM. A power glitch tripped off the Booster West Anode Power Supply and a Pbar Accumulator RF station (ARF1); they both reset. EE Support personnel reported at 6:47 PM that they brought a Switchyard quadrupole power supply (S:QP3) online. Operations established beam to the Meson FTBF T-1008 experiment at 7:01 PM. Operators put the TEV through a low beta magnets squeeze at 7:52 PM. Operations disabled beam to T-1008 at 10:01 PM. Operations put the TEV into shot setup at 10:04 PM. At 10:29 PM, the newly installed 30-ton air condition (AC) unit for MiniBooNE at the MI-12 service building tripped off on a ground fault indication. Operations disabled MiniBooNE and FESS personnel and operators installed two portable AC units. At 10:50 PM, the MI Power Supply Interlock Loop tripped off on a ground fault indication; it reset. At 11:06 PM, the TEV quenched during collimating; operators and TEV personnel investigated. At 11:53 PM, operators reported that all of the TEV RF stations tripped off. Saturday July 23 The midnight shift began with Operations monitoring the stash (59.6E10) and the Pbar stack (11.7E10), and with MiniBooNE down due to AC problems in MI-12. During the midnight shift, Operations disabled beam to T-1008 at 12:06 AM. Operations established beam to MiniBooNE at 1:01 AM. At 2:23 AM, a lightning strike brought down the MI abort kickers, the MI beam permit, and the Pbar beam permit; everything reset. The strike also took out the CHL nitrogen plant; Cryo System technicians put the plant into recovery mode. At 2:28 AM, a lightning strike took out the Recycler stash and several Booster RF stations. At 12:54 AM, lightning took down the MI RF system; it reset. At 3:49 AM, Casey’s Pond ICW pumps lost power; the High Voltage Coordinator came in to help with the power outages. At 4:01 AM, the P2 transfer line beam permit tripped off due to a quadrupole power supply (I:QF12) trip. At 4:11 AM, power had been restored to Casey’s pond pump house, but FESS personnel reported that the #1 pump had a fault; they got the #2 pump on. At 5:09 AM, the High Voltage Coordinator reported that Feeder 32 was down; he and FESS personnel had it powered back up by 5:31 AM. Operators put the TEV through a dry squeeze at 6:45 AM. Operations established beam to T-1008 at 7:11 AM, after fixing some problems. At 7:55 AM, operators put the TEV through a dry squeeze, but during the squeeze, the leads at TEV sectors A4 and E3 timed out and caused another quench. Operators investigated and discovered that while they had tried to resolve some script problems the TEV had stopped at 90 GeV and remained at 90 GeV during the ramp, acceleration, and low beta squeeze. During the day shift, Controls personnel came in to work on crates and other issues. Operators reported at 9:31 AM, that the C4 cryo systems were recovering. BRF14 tripped off; it reset. Operators restarted the MiniBooNE auto-tune program. A FESS person reported at 11:08 AM that he needed to get into MI-12 for AC investigations; operators turned off the MiniBooNE beam for about 40 minutes. BRF12 and 13 tripped off; they reset. Operators put the TEV through a dry squeeze at 1:56 PM and through a wet squeeze at 2:47 PM. Operations put the TEV into shot setup at 3:55 PM. During the evening shift, Operations established store 8923 at 5:26 PM with an average initial luminosity (AIL) of 309.78E30. Operators disabled beam to MiniBooNE at 11 PM due to high temperatures. Operations disabled beam to T-1008 at 11:32 PM. At 11:35 PM, a lightening strike caused store 8923 to abort; there was no quench. Many other systems tripped off. Operators reported at 11:49 PM that two Booster quadrupole power supplies (B:Q3 and B:Q10) had no output; operators called in a Booster person. Sunday July 24 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE off due to high temperatures. During the midnight shift, operators reported at 1:46 AM, that while they were putting the TEV through 980 GeV ramps, they discovered problems with the A1 leads; operators reported finding a problem with input/output board in the A1 CAMAC crate; a Cryo System tech came in to replace the board. At 2:25 AM, BRF13 tripped off for the first of eight times in about five hours; it reset. Operators put the TEV through dry and wet squeezes. At 6:51 AM, Operations consulted with the Run Coordinator and decided to delay shot setup until a nearing storm passed by. During the day shift, operators prepared MI-12B for access at 8:17 AM to check magnets and sump pumps. Operations put the TEV into shot setup at 9:11 AM and established store 8926 at 10:20 AM with an AIL of 350.08E30. Operations established beam to T-1008 at 10:55 AM. Water Works personnel reported from MI-12B at 11:47 AM that quadrupole magnet 874 was clogged; later they reported that the magnet wasn’t clogged, but the hoses to the magnet were. Water Works personnel searched for the right sized hoses. During the evening shift, Water Works personnel completed their Q874 repairs at 6:25 PM. Operations established beam to MiniBooNE at 6:29 PM. Monday July 25 The midnight shift began with Operations monitoring store 8926 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations disabled beam to T-1008 at midnight. At 12:30 AM, BRF1 and other Booster equipment (Noker and the Spellman HV) tripped off; a Booster person came in to investigate. Operations terminated store 8926 at 2:07 AM, put the TEV into shot setup, and established store 8927 at 3:18 AM with and AIL of 386.1E30. The Plans for Monday The plans for today are to stack, stash and store. The Run Coordinator will allow MiniBooNE to turn off and access to repair their AC at MI-12. The NuMI V108 magnet needs new cables pulled from the power supply down to the magnet. Preparations for this will start today. MI will turn off on Tuesday morning (7/26/11) at 6 AM for about 12 hours for the cable pull. |
|
More Information Comments and Suggestions |
| last modified 07/25/2011 email Fermilab |
| Security, Privacy, Legal | |